OpenStudy (anonymous):

two numbers sum 48 and have a ratio of 2 to 6. which are there numbers? anwer using two numbers (e.g.: 1 and 2)

OpenStudy (anonymous):

2x+6x=48 8x=48 x=48/8 x=6 numbers are 2(6)=12 and 6(6)=36

OpenStudy (anonymous):

Nice job math, i was all like Fk that lol